Suppose a 0.270 -kg mass on a spring-loaded gun that has been compressed 0.210 m has elastic potential energy of 24.0 J. How high above the spring's equilibrium point can the gun fire the mass if the gun is fired straight up?
Answer #1

From the conservation of energy

Initial elastic potential energy = Final gravitational potential energy

there fore 24 J = mgh

height above the equilibrium point h = 24/mg = 24/(.270*9.8) = 9.07 meters